OpenStudy (lordhelix8th):

WILL MEDAL AND FAN What is the exact area of a circle whose diameter is 14 cm? A. 14 straight pi cm2 B. 28 straight pi cm2 C. 49 straight pi cm2 D. 196 straight pi cm2

OpenStudy (lordhelix8th):

@ineedhelpeeps @starlord6200 @SydtheKid913

OpenStudy (anonymous):

i think that its c

OpenStudy (sydthekid913):

A= pi * r^2 plug in your values

OpenStudy (lordhelix8th):

d?

OpenStudy (lordhelix8th):

thats what i came closest too am i right

OpenStudy (sydthekid913):

Radius is the diameter divided by 2. 14/2=7; 7^2=49; 49*pi=? Yes I believe you are correct

OpenStudy (lordhelix8th):

ok

OpenStudy (lordhelix8th):

let me round just in case thx

OpenStudy (sydthekid913):

This is an unusual question... My answer doesn't match at all. What did you get?

OpenStudy (lordhelix8th):

i got 153.84

OpenStudy (sydthekid913):

So did I

OpenStudy (lordhelix8th):

hmmmmmmmm

OpenStudy (sydthekid913):

Hmmmm

OpenStudy (sydthekid913):

I'd go for d... I wish I could give you a more exact answer but I'm not sure what they're looking for.

OpenStudy (lordhelix8th):

ok @arabpride @Daniellelovee @gardenGnostic @Awolflover1

OpenStudy (sydthekid913):

@helpblahblahblah

OpenStudy (lordhelix8th):

@dan815 @I_love_my_nieces @Wolf19 @Samara8954

OpenStudy (lordhelix8th):

@King.Void. @kidrah69 @khalilforthewin @Kymstokes @KodyKthnx @DARTHVADER2900

OpenStudy (lordhelix8th):

@DarkMoonZ

OpenStudy (anonymous):

Ok, so how do you find the area of a circle? Do you know?

OpenStudy (lordhelix8th):

A= pi * r^2

OpenStudy (anonymous):

ok, so then you'd use that equation with 14 cm

OpenStudy (lordhelix8th):

i got 153.94

OpenStudy (lordhelix8th):

hello

OpenStudy (lordhelix8th):

@AJ01

OpenStudy (anonymous):

area = pi * r^2 =pi * (14/2)^2 = pi * 7^2= 49pi

OpenStudy (lordhelix8th):

u sure

OpenStudy (anonymous):

the answer is c

OpenStudy (anonymous):

100%

OpenStudy (lordhelix8th):

ok

OpenStudy (lordhelix8th):

Caley has a bag that contains 9 blue marbles and 13 yellow marbles. She selects a marble at random, and then, without replacing the first one, selects another marble at random. What is the probability that Caley selects a blue marble and then a yellow marble? Round your answer to the nearest percent. P(blue and yellow) ≈ %

OpenStudy (lordhelix8th):

can u help with 1 more @AJ01

OpenStudy (anonymous):

lol i was right

OpenStudy (lordhelix8th):

lol

OpenStudy (lordhelix8th):

@starlord6200 CAN U HELP WITH THIS

OpenStudy (anonymous):

idk thid